# DualDynamics: Synergizing Implicit and Explicit Methods for Robust Irregular Time Series Analysis

### v4

@@ Introduction
We are interested in computing the trace of $\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}s}}$ in Eq. (13), which is a key quantity to understand the change in log-density. We use Hutchinson’s trace estimator111 Hutchinson’s trace estimator states that for a square matrix $A$, the trace of $A$ can be estimated as: $\text{Tr}(A)\approx\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}A\mathbf{v}],$ where $\mathbf{v}$ is a random vector with each element drawn independently from a distribution with zero mean and unit variance. (Hutchinson 1989) used to estimate the trace of a matrix efficiently, especially in situations where the matrix is large and computing the trace directly is computationally expensive [^2] Hutchinson’s trace estimator significantly reduces computational complexity from $\mathcal{O}({d_{z}}^{2})$ for direct trace computation to $\mathcal{O}(d_{z})$ for matrix-vector multiplications. (Adams et al. 2018; Han, Malioutov, and Shin 2015; Grathwohl et al. 2018).
@@
$\text{Tr}\left(\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\right)\approx\mathbb{E}_{\mathbf{v}}\left[\mathbf{v}^{\top}\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\mathbf{v}\right].$ (15)
@@
$\qquad\approx-\mathbb{E}_{\mathbf{v}}\left[\int_{0}^{t}\mathbf{v}^{\top}\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\mathbf{v}\,{\textnormal{d}}s\right]$
$\qquad=-\mathbb{E}_{\mathbf{v}}\left[\int_{0}^{t}\mathbf{v}^{\top}\frac{\partial\bm{J}_{{\mathcal{G}}}(\hat{{\bm{z}}}(t))}{\partial t}\mathbf{v}\,{\textnormal{d}}s\right].$ (16)
@@
Table 1: Average classification performance on 18 datasets under regular and three missing rates (Values in parentheses show the average of 18 standard deviations. The top-ranking and runner-up choices are emphasized.)
@@
Figure 4: Forecasting performance versus the number of parameters on MuJoCo dataset with regular scenario
@@
Figure 5: Forecasting performance versus the number of parameters on Google dataset with regular scenario
@@ Discussion
Figure 6: Ablation study of different flow model configurations versus conventional MLP
@@ Conclusion
Reproducibility Statement
We make our code and implementation details publicly accessible for reproducibility purposes. Code is available at https://github.com/yongkyung-oh/DualDynamics.
@@ Appendix A Experimental Details
6:    Compute cross-entropy loss $\mathcal{L}$ using $\hat{\mathbf{z}}(T)$ and target $y$.
$\mathcal{L}=-\sum_{i=1}^{C}y_{i}\log(\text{softmax}(MLP(\hat{\mathbf{z}}(T);\theta_{MLP}))_{i})$
@@ Appendix B Detailed Results for “Robustness to Dataset Shift”
We conducted classification experiments across 18 diverse datasets, in three distinct domains, from the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository  [^3] http://www.timeseriesclassification.com/ (Bagnall et al. 2018) using the python library sktime (Löning et al. 2019). As outlined in Table 7, the datasets in the Sensor domain are univariate, whereas the others encompass a mix of univariate and multivariate data. Addressing the challenge posed by varying lengths of time series, the strategy of uniform scaling was employed to align all series to the dimension of the longest series (Keogh 2003; Yankov et al. 2007; Gao and Lin 2018; Tan et al. 2019). The partitioning of the dataset into training, validation, and testing subsets was executed in proportions of 0.70, 0.15, and 0.15 respectively. Subsequent to this partitioning, a random assortment of missing observations was introduced for each variable. Lastly, these modified variables are integrated into combined dataset.
@@
Table 11: Ablation study of the proposed method on 18 datasets under regular and three missing rates
(Values in parentheses show average of 18 standard deviations)
@@ Appendix E Detailed Results for “Forecasting with the Partial Observations”
Table 20: Results of hyperparameter tuning for PhysioNet Sepsis Dataset with observation intensity
(Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
@@
Table 21: Results of hyperparameter tuning for PhysioNet Sepsis Dataset without observation intensity
(Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
@@
Table 22: Results of hyperparameter tuning for MuJoCo Dataset
(Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
@@
Table 23: Results of hyperparameter tuning for Google Dataset
(Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)

### v6

@@ Introduction
We are interested in computing the trace of $\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}s}}$ in Eq. (13), which is a key quantity to understand the change in log-density. We use Hutchinson’s trace estimator111 Hutchinson’s trace estimator states that for a square matrix $A$, the trace of $A$ can be estimated as: $\text{Tr}(A)\approx\mathbb{E}_{\bm{v}}[\bm{v}^{\top}A\bm{v}],$ where $\bm{v}$ is a random vector with each element drawn independently from a distribution with zero mean and unit variance. (Hutchinson 1989) used to estimate the trace of a matrix efficiently, especially in situations where the matrix is large and computing the trace directly is computationally expensive [^2] Hutchinson’s trace estimator significantly reduces computational complexity from $\mathcal{O}({d_{z}}^{2})$ for direct trace computation to $\mathcal{O}(d_{z})$ for matrix-vector multiplications. (Adams et al. 2018; Han, Malioutov, and Shin 2015; Grathwohl et al. 2018).
@@
$\text{Tr}\left(\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\right)\approx\mathbb{E}_{\bm{v}}\left[\bm{v}^{\top}\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\bm{v}\right].$ (15)
@@
$\qquad\approx-\mathbb{E}_{\bm{v}}\left[\int_{0}^{t}\bm{v}^{\top}\bm{J}_{\frac{\mathrm{d}{\mathcal{G}}}{\mathrm{d}t}}(\hat{{\bm{z}}}(t))\bm{v}\,{\textnormal{d}}s\right]$
$\qquad=-\mathbb{E}_{\bm{v}}\left[\int_{0}^{t}\bm{v}^{\top}\frac{\partial\bm{J}_{{\mathcal{G}}}(\hat{{\bm{z}}}(t))}{\partial t}\bm{v}\,{\textnormal{d}}s\right].$ (16)
@@
Table 1: Average classification performance on 18 datasets under regular and three missing rates (Values in parentheses show the average of 18 standard deviations. The top-ranking and runner-up methods are emphasized.)
@@
Figure 4: Forecasting performance versus the number of parameters on MuJoCo dataset with regular scenario (different flow model (with remarked *) versus conventional MLP)
@@
Figure 5: Forecasting performance versus the number of parameters on Google dataset with regular scenario (different flow model (with remarked *) versus conventional MLP)
@@ Discussion
Figure 6: Ablation study of different flow model (with remarked *) versus conventional MLP
@@ Appendix A Experimental Details
6:    Compute cross-entropy loss $\mathcal{L}$ using $\hat{\bm{z}}(T)$ and target $y$.
$\mathcal{L}=-\sum_{i=1}^{C}y_{i}\log(\text{softmax}(MLP(\hat{\bm{z}}(T);\theta_{MLP}))_{i})$
@@ Appendix B Detailed Results for “Robustness to Dataset Shift”
We conducted classification experiments across 18 diverse datasets, in three distinct domains, from the University of East Anglia (UEA) and the University of California Riverside (UCR) Time Series Classification Repository  [^3] http://www.timeseriesclassification.com/ (Bagnall et al. 2018; Dau et al. 2019) using the python library sktime (Löning et al. 2019), as outlined in Table 7. Addressing the challenge posed by varying lengths of time series, the strategy of uniform scaling was employed to align all series to the dimension of the longest series (Keogh 2003; Yankov et al. 2007; Gao and Lin 2018; Tan et al. 2019). The partitioning of the dataset into training, validation, and testing subsets was executed in proportions of 0.70, 0.15, and 0.15 respectively. Subsequent to this partitioning, a random assortment of missing observations was introduced for each variable. Lastly, these modified variables are integrated into combined dataset.
@@
Table 11: Ablation study of the proposed method on 18 datasets under regular and three missing rates (Values in parentheses show average of 18 standard deviations)
@@ Appendix E Detailed Results for “Forecasting with the Partial Observations”
Table 20: Results of hyperparameter tuning for PhysioNet Sepsis Dataset with OI (Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
@@
Table 21: Results of hyperparameter tuning for PhysioNet Sepsis Dataset without OI (Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
@@
Table 22: Results of hyperparameter tuning for MuJoCo Dataset (Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)
@@
Table 23: Results of hyperparameter tuning for Google Dataset (Memory usage in MB, and the number of parameter in ×$\times 10^{3}$ unit.)